# E. Gover11footnotetext: Corresponding author.    D. Veske    M.B. Demirkoz

### v1

@@ 7 Neutron Beam Distribution Characteristics
where the fitted parameters were extracted from ROOT file. The calculated Gaussian parameters were as follows: A=749.252, $y_{0}$=0.009, $z_{0}$=2.942 with standard deviations $\sigma_{y}$=14.874, $\sigma_{z}$=14.220. For comparison, the FWHM of incoming proton beam was 1.5 cm with $\sigma=0.636$. Figure 16 and 17 shows the LEGO plot of beam profile in 2D and 3D respectively. The neutron beam projection spans a large area, which extends to $\sim 100\text{ cm}$ in each direction. To better collimate the beam and obtain better focused neutron distribution, novel collimator designs should be implemented.
@@ 8 Conclusion
This work presented the preliminary trial run of a proposed proton-beryllium irradiation experiment. Different nuclear simulation toolkits may present different results, hence this study was conducted to address the differences between FLUKA and Geant4. All simulation results in this work should be cross-checked with experimental beryllium irradiation data from EXFOR[15]. The numerical data obtained from FLUKA using the PRECISION physics as default setting is in more parallel with Geant4 when the physics list specified in Section 3.2 is registered.
Beryllium is a compelling target material for accelerator-driven neutron sources. The significant flexibility in tailoring the neutron energy spectrum (which can be done by varying the incident particle energy in the accelerator and the moderators) enables the feasibility of a wide range of neutron-based studies. Moreover, the isotropic nature of the produced neutron distribution is dependent on both the target orientation, and the kinetic energy of the incoming projectile. This angular aspect provides an additional degree of freedom for shaping the neutron beam. The low atomic number of beryllium (Z=4) and relatively low neutron binding energy (2.4 MeV) enables efficient neutron production through reactions such as ${}^{9}\mathrm{Be}(p,n)^{9}\mathrm{B}$ and ${}^{9}\mathrm{Be}(d,n)^{10}\mathrm{B}$. Its relatively high melting point and thermal conductivity makes it convenient to implement liquid water as primary cooling mechanism, which would be an efficient and cheap alternative to cryogenic cooling. However, beryllium should be handled with caution, as it is highly toxic, volatile under high-intensity irradiation and requires strict safety protocols.
@@ Acknowledgments
The author expresses gratitude to Dr. Doğa Veske and Prof. Dr. M. Bilge Demirköz for their valuable support throughout this work.

### v2

@@ 5.3 Neutron Beam Distribution Characteristics
where the fitted parameters were extracted from ROOT file. The calculated Gaussian parameters were as follows: A=749.252, $y_{0}$=0.009 cm, $z_{0}$=2.942 cm with standard deviations $\sigma_{y}$=14.874 cm, $\sigma_{z}$=14.220 cm. For comparison, the FWHM of incoming proton beam was 1.5 cm with $\sigma=0.636$ cm. Figure 16 and 17 shows the LEGO plot of beam profile in 2D and 3D respectively. The neutron beam projection spans a large area, which extends to $\sim 100\text{ cm}$ in each direction.
@@ 6 Conclusion
This work presented the preliminary trial run of a proposed proton-beryllium irradiation experiment. The numerical data obtained from FLUKA using the PRECISION physics as default setting is in more parallel with Geant4 when the physics list specified in Section 3.1 is registered. We found at higher neutron energies, FLUKA seems to score higher neutron fluence than Geant4, while at lower energies these two simulation toolkits agree on fluence estimation.
Beryllium is a compelling target material for accelerator-driven neutron sources. The significant flexibility in tailoring the neutron energy spectrum (which can be done by varying the incident particle energy in the accelerator and the moderators) enables the feasibility of a wide range of neutron-based studies. The angular aspect provides an additional degree of freedom for shaping the neutron beam. The low atomic number of beryllium (Z=4) and relatively low neutron separation energy (1.67 MeV) enables efficient neutron production through reactions such as $^{9}\mathrm{Be}(p,n)^{9}\mathrm{B}$ and $^{9}\mathrm{Be}(d,n)^{10}\mathrm{B}$. Its relatively high melting point and thermal conductivity makes it convenient to implement liquid water as primary cooling mechanism, which would be an efficient and cheap alternative to cryogenic cooling. However, beryllium should be handled with caution, as it is highly toxic, volatile under high-intensity irradiation and requires strict safety protocols.
@@ Acknowledgments
DV is grateful to The Scientific and Technological Research Council of Türkiye for their support through the grants 123C484 and 125F379. DV acknowledges further support from the European Union’s Horizon Europe Research and Innovation Programme under the Marie Skłodowska-Curie Actions (MSCA) COFUND Programme with the grant number 101081645 and The Scientific and Technological Research Council of Türkiye with the grant number 123C213. We extend our sincere gratitude to Middle East Technical University for their generous support.